### v1

# Conceptualization of this study, Methodology, Software 2]organization=Academy of Mathematics and Systems Science, Chinese Academy of Sciences, city=Beijing, postcode=100190, country=China

Abstract
Recent advances on time series forecasting mainly focus on improving the forecasting models themselves. However, managing the length of the input data can also significantly enhance prediction performance. In this paper, we introduce a novel approach called Optimal Starting Point Time Series Forecast (OSP-TSP) to capture the intrinsic characteristics of time series data. By adjusting the sequence length via leveraging the XGBoost and LightGBM models, the proposed approach can determine optimal starting point (OSP) of the time series and thus enhance the prediction performances. The performances of the OSP-TSP approach are then evaluated across various frequencies on the M4 dataset and other real-world datasets. Empirical results indicate that predictions based on the OSP-TSP approach consistently outperform those using the complete dataset. Moreover, recognizing the necessity of sufficient data to effectively train models for OSP identification, we further propose targeted solutions to address the issue of data insufficiency.
keywords:
[…]
1 Introduction
Time series forecasting plays a crucial role in practical applications across various fields including services, tourism, finance, meteorology and many others. Before the rise of machine learning — particularly deep learning — traditional forecasting methods predominantly relied on statistical models like Exponential Smoothing (ETS) Hyndman et al. (2008) and ARIMA Box et al. (2015). These models, which use past observations to construct linear functions for predicting future trends, have been widely employed in forecasting tasks for decades. In recent years, deep learning techniques, such as Recurrent Neural Networks (RNNs) Lipton (2015) and ConvTrans Li et al. (2019), have gained popularity in time series forecasting. These approaches excel at capturing complex nonlinear patterns in the data, leading to more accurate predictions.
Previous literature on time series modeling has mainly concentrated on model selection and optimization, such as enhancing forecast accuracy through the integration of forecasting, ensemble learning, and artificial neural networks. These techniques rely on learning from the entire dataset to make predictions. Typically, providing a predictive model with all available data allows it to form a comprehensive understanding of historical patterns, thereby enhancing the model’s predictive capabilities. With recent advances in data storage technologies, we can now easily store and retrieve vast amounts of time series data for model training. However, when structural changes occur in the data, using the full dataset may lead to predictions that place too much weight on long-term trends and potentially overlooking recent shifts that have emerged after these changes.
For example, the flow of tourists between China and Japan had been steadily increasing since 1979, with a significant surge after 2010. However, the outbreak of COVID-19 in 2020 had a profound global impact, and the subsequent quarantine measures imposed by various countries caused a sharp decline in travel. As a result, the number of Chinese tourists visiting Japan in 2020 and 2021 plummeted to levels reminiscent of the pre-2010 era. Currently, tourism is in a phase of recovery, gradually returning to the levels observed before 2019. If we use total tourist numbers data for future predictions, it might overly emphasize the consistent growth observed before 2019, failing to account for the tourism industry’s recovery from the pandemic. This could lead to significant inaccuracies in the forecast. While certain time series models, such as ETS and LSTM, give more weight to recent data to minimize the noise from long-term trends, it may be more appropriate in this case to exclude data prior to the COVID-19 outbreak to improve forecast accuracy. However, simply truncating data may not always result in the most optimal forecast. In fact, we should expect that there exists a specific point in the time series such that if we begin the forecast from that point, the forecast error can be minimized. We refer to this as the optimal starting point (OSP) of the time series. In the previous example, we subjectively chose the starting point of a new time series based on the onset of the COVID-19 pandemic. However, for other time series data, we may not have clear prior knowledge, as we do with tourism data, to pinpoint the location of structural changes that would inform data truncation. Moreover, there are various ways to truncate a time series, making it difficult to identify this point subjectively. Therefore, our goal is to develop a machine learning model that can help us automatically determine the OSP of a time series. This is the primary focus of this article.
It is worth noting that different time series are likely to vary in length, which brings challenge for model training. When feeding a set of time series into a model, the length inconsistency can complicate the training process. To address this issue, we consider an approach similar to change point detection, where we extract consistent features from each time series before putting them into the model for training. These features, such as length, frequency, and other relevant metrics, provide important information into the characteristics of the time series and play a crucial role in improving forecasting accuracy. In existing research, time series features are often directly input into prediction models to enhance accuracy. For instance, models like LightGBM Mamonov et al. (2022) have been successfully applied to time series forecasting. Additionally, time series features are also used in meta-learning for combinatorial forecasting, where they help assign weights to different base models. For example, the FFORMA (Feature-based Forecast Model Averaging) method, introduced by Montero-Manso et al. (2021), uses 42 statistical features (implemented via the R package tsfeatures) to estimate optimal weights for nine traditional models, trained using XGBoost. In this study, we will use LightGBM and XGBoost models learn the characteristics of the time series and determine the location of the optimal starting point.
The rest of the paper is organized as follows. Section 2 introduces the related works from optimal starting point determination and time series feature extraction. Section 3 introduces the base prediction model and method for improving predictions. Section 4 demonstrates the performances of the proposed method through empirical studies on M4 data and other datasets. Finally, Section 5 conclude this article with some discussions.
2 Related Work
2.1 Optimal starting point determination
The determination of the optimal starting point of time series is similar to the change point detection. Change point detection algorithms are a suite of methods designed to identify when significant shifts occur in time series data. The existing literature of change point detection can be categorized into supervised and unsupervised approaches. The supervised approaches include decision trees Zheng et al. (2008), hidden Markov models, Gaussian mixture models Han et al. (2012), and many others. These models are frequently employed in transportation domains, such as using accelerometer or audio data to detect changes in human activity states (e.g., walking versus running). Unsupervised learning approaches, on the other hand, do not require training data. Instead, they detect patterns and changes directly from the time series data. These methods are suitable for unlabeled data and can be used without prior knowledge. These methods typically detect changes by computing or estimating a particular characteristic value of time intervals. For instance, Kawahara and Sugiyama (2011) proposes directly estimating the ratio of probability densities and calculating the likelihood ratio between reference and test intervals to detect changes. Similarly, Liu et al. (2013) introduces a novel statistical change point detection algorithm based on relative Pearson divergence, which accurately and efficiently estimates the non-parametric divergence between time series samples using direct density ratio estimation. Additionally, a change point detection algorithm based on subspace identification has been proposed, leveraging the approximate equivalence between the column space of the observability matrix and the space spanned by sub-sequences of the time series data Kawahara et al. (2008). This approach evaluates incoming sub-sequences of data to achieve effective change point detection.
2.2 Time series feature extraction
Traditional time series feature extraction approaches are based on statistical methods, such as summary statistics (e.g., mean, variance, skewness, and kurtosis) Hamilton (1994).These features have expanded over time to address more complex needs. Statistical methods are simple, intuitive, and easy to implement. With the growing demand for time series analysis, several software packages have emerged for automated feature extraction. For instance, Christ et al. (2017) introduced the tsfresh open-source library in Python, which implements 63 feature extraction methods, including statistical, correlation, and stationarity features. It is based on extensible hypothesis testing (like the FRESH algorithm) and automatically identifies and extracts the most relevant features for tasks such as classification or regression, making it scalable and suitable for large-scale time series data. Similarly, Barandas et al. (2020) introduced TSFEL, a Python package offering over 60 feature extraction methods across time, statistical, and frequency domains. TSFEL allows customization through an online interface or directly as a Python package, making it ideal for rapid exploratory data analysis. Additionally, it provides an evaluation of computational complexity, helping users estimate the computational cost of feature extraction early in the machine learning pipeline.
There are also other ways to construct time series features. Mörchen (2003) proposes a novel method for time series feature selection based on DWT and DFT, which outperforms traditional methods in terms of energy preservation. This method has the potential to improve clustering results and reduce the size of classification rule sets. Zhang et al. (2006) presents an unsupervised feature extraction method based on orthogonal wavelet transform. This method automatically selects the best wavelet decomposition scale to extract key features from time series, with a time complexity of $O(mn)$, where $m$ is the number of time series and $n$ is the length of each time series. Experimental results show that the features extracted by this method can improve the quality of time series clustering. Olszewski (2001) proposes a generalized feature extraction method for structural pattern recognition in time series data. Based on six types of modulation commonly used in signal processing, this method designs six structural detectors to identify these fundamental structures in time series. In order to analyze which features will affect the determination of the optimal starting point and improve the interpretability of the model, we choose to use the traditional method of constructing features.
3 Methodology
The Optimal Starting Point Time Series Forecast (OSP-TSP) algorithm contains two main steps. First, we train a supervised model (OSP) to compute the optimal starting point of the time series. Second, the optimal starting point is then used to truncate the new sequence data for the final prediction model. The key idea of the OSP-TSP algorithm is to utilize machine learning techniques to identify key characteristics of a time series, thereby predicting the interval where the optimal starting point lies. These predictions are then used to improve the accuracy of time series forecasts. Essentially, this model can be applied to any time series data to enhance prediction precision. For example, it can be employed on simulated time series generated by methods like GRATIS Kang et al. (2019). Next, we will provide a detailed introduction to the proposed Algorithm.
3.1 Simplified Interval Forecast
We define the optimal starting point, denoted as $y_{bst}$, of a time series in the following manner. Given a complete time series dataset $\{y_{1},y_{2},...,y_{T}\}$, our objective is to generate a prediction for the subsequent $h$ periods that minimizes the target forecast error metrics (such as mean absolute scaled error, root mean squared error, etc.). We randomly select a point, $y_{st}$, from the sequence to serve as the starting point of a new sequence. This new sequence, ${y_{st},y_{st+1},...,y_{T}}$, is formed by truncating the original data. We then feed this new sequence into the time series prediction model $M$, forecast $h$ periods into the future, and compute the final target error. The optimal starting point, $y_{bst}$, is the point that yields the smallest target error among all possible $y_{st}$. However, it’s worth noting that accurately identifying the optimal starting point of a time series is a laborious task. In the absence of any additional information, for a sequence of length $T$, we would need to construct $T-1$ forecast models and compare the resulting errors to determine the starting point that minimizes the forecast error. If our objective is to accurately predict the position of $y_{bst}$, we would need to continuously execute the aforementioned operations during the model training process to assign labels to the training set data. This precise method, however, is time-consuming and exhibits a certain degree of instability.
To mitigate the complexity of training the model while maintaining a balance between accuracy and computational cost, we have simplified the prediction goal. Instead of predicting the exact optimal starting point, we aim to predict the optimal forecast starting interval. This can be intuitively understood as the time series interval in which the optimal starting point $y_{bst}$ is located. During the model training phase, we divide each training data sequence evenly into $m$ sub-intervals. For each time series, we only calculate and store the model target prediction error starting from $n$ points between equal partitions in each sub-interval, and subsequently mark a certain interval as the location of the optimal starting point.
For each original time series, we construct $mn$ sub-time series and $mn$ sets of prediction errors for future $h$ periods. From these prediction errors, we can identify the interval $y$ where the actual optimal starting point is located, which can then be used as the label for subsequent machine learning tasks.For each time series, after calculating the target forecast error of $mn$ series, there are two methods for marking the interval where the optimal starting point of prediction is, as follows:
  1. [1)]
  2. 1.
    For $m$ groups of intervals, we compare the $n$ sets of errors formed by $n$ points in each sub-interval, and use the minimum value of target forecast error as the error value of that sub-interval. Then we choose the smallest value as the interval y where the optimal starting point is located.
  3. 2.
    For $m$ groups of intervals, calculate the average error formed by $n$ points in each sub-interval, and take the average error as the error value of that sub-interval. Then we choose the smallest value as the interval $y$ where the optimal starting point is located.
The above two different methods of determining the interval where the actual optimal starting point is located will affect the subsequent problem of forecast accuracy, and we will analyze this effect in the empirical analysis..
3.2 Feature Extraction
The R package tsfeatures provides methods to extract features from time series data, and the specific variable names and corresponding meaning descriptions are shown in Table 1. We use the tsfeatures function for each time series in the M4 data to obtain the corresponding feature data Yang and Hyndman (2022).
Table 1: Number of M4 series per data frequency and domain
Variable Description
frequency Data frequency(1 for annual data, 4 for quarterly and 12 for monthly)
nperiods Set to 1 for non-seasonal data.
seasonal_periodis The number of seasonal periods in the data (determined by the frequency of observation, not the observations themselves).
trend Strength of trend of the the STL decomposition approach.
spike The spikiness of a time series and is computed as the variance of the leave-one-out variances of the remainder component.
linearity The linearity of a time series calculated based on the coefficients of an orthogonal quadratic regression.
curvature The curvature of a time series calculated based on the coefficients of an orthogonal quadratic regression.
e_acf1 The first autocorrelation coefficient of the the autocorrelation function of $e_{t}$.
e_acf10 The sum of the first ten squared autocorrelation coefficients of the the autocorrelation function of $e_{t}$.
seasonal_strength Strength of seasonality of the the STL decomposition approach.
peak The size and location of the peaks in the seasonal component.
trough The size and location of the troughs in the seasonal component.
entropy The forecastability of a time series, where low values indicate a high signal-to-noise ratio, and large values occur when a series is difficult to forecast.
x_acf1 The first autocorrelation coefficient of the the autocorrelation function of the series.
x_acf10 The sum of squares of the first ten autocorrelation coefficient of the the autocorrelation function of the series.
diff1_acf1 The first autocorrelation coefficient of the autocorrelation function of the differenced series.
diff1_acf10 The sum of squares of the first ten autocorrelation coefficient of the the autocorrelation function of the differenced series.
diff2_acf1 The first autocorrelation coefficient of the autocorrelation function of the twice-differenced series.
diff2_acf10 The sum of squares of the first ten autocorrelation coefficient of the the autocorrelation function of the twiced-differenced series.
seas_acf1 The first autocorrelation coefficient of the autocorrelation function of the seasonality.
3.3 Training the Optimal Starting Point Model
Upon extracting the features from all training data, we feed them into the Optimal Starting Point (OSP) model $C$ as independent variables. The dependent variable $y$ represents the interval in which the optimal starting point is located. The OSP model $C$ chosen for this study primarily relies on machine learning algorithms, such as XGBoost or LightGBM.
  • XGBoost XGBoost is an efficient implementation of the Gradient Boosting Decision Tree (GBDT) and offers improvements over the original GBDT Chen and Guestrin (2016). As a forward addition model, its core strategy is to ensemble multiple weak learners into a strong one using boosting. The output of each decision tree is the difference between the target value and the prediction result of all preceding trees. The final result is obtained by summing up all these outputs:
    $\widehat{y}_{i}^{(t)}=\sum_{k=1}^{t}f_{k}(x_{i})=\widehat{y}_{i}^{(t-1)}+f_{t}(x_{i}),$
    where $y_{i}^{(t)}$ is the prediction result for sample i after the $t$th, $y_{i}^{(t-1)}$ is the prediction results for the perecious $t-1$ trees and $f_{t}(x_{i})$ is the $t$th DT.
    XGBoost is an efficient implementation of GBDT in that it successfully alleviates the overfitting problem by introducing a regularization function $\Omega(f)$:
    $\Omega(f)=\gamma T+\frac{1}{2}\lambda\sum_{t=1}^{T}t^{2},$
    where $\gamma$ and $\lambda$ are hyper-parameters used to control the complexity, and $T$ is the number of lead nodes.
    Therefore, the objective function of XGBoost based on the GBDT optimization objective can be expressed as:
    $\min\left[L\{F_{t}(x),y\}+\Omega(f)+\zeta\right]=\min\left\{\sum_{i=1}^{N}\Big{(}\widehat{y}_{i}^{(t)},y_{i}^{(t)}\Big{)}+\sum_{t=1}^{T}\Omega(f_{t})+\zeta\right\}.$
    In this manner, we can iteratively optimize the loss function to achieve the sum of all Decision Tree (DT) predictions in XGBoost through the application of an additive boosting-type ensemble model.
  • LightGBM
    LightGBM is a framework that implements the Gradient Boosting Decision Tree (GBDT) algorithm. It supports efficient parallel training and offers advantages such as faster training speed, lower memory consumption, improved accuracy, and distributed support (Ke et al. (2017)). The primary motivation behind the development of LightGBM was to address the challenges faced by GBDT in handling massive data. It aims to accelerate the training of GBDT models without compromising accuracy, thereby enabling GBDT to be used more effectively and swiftly in industrial applications.
    Initially, LightGBM introduces a leaf-wise growth strategy with a depth limit. This strategy avoids indiscriminate treatment of leaves at the same level, which can reduce more errors and achieve better accuracy with the same number of splits, while avoiding unnecessary overhead.
    Furthermore, the Gradient-based One-Side Sampling (GOSS) algorithm adopts a sample reduction approach. It excludes most samples with small gradients and calculates the information gain using only the remaining samples. This approach provides a balanced algorithm in terms of reducing the amount of data and ensuring accuracy.
    High-dimensional data tend to be sparse, and this sparsity has inspired the design of a lossless method to reduce the dimensionality of features. As such, the Exclusive Feature Bundling (EFB) algorithm posits that the number of features can be reduced if certain features are fused and bound together.
  1. [1)]
  2. 1.
    The dependent variable is qualitative data, with a total of $m$ categories, and the categorical prediction model is constructed.
  3. 2.
    Considering that the interval where the optimal starting point is located is continuous on the time axis, the dependent variable is regarded as quantitative data and thus a regression prediction model should be constructed. Since the direct prediction results are often not integers, the closest integer on the number axis which between 1 and $m$ is taken as the interval where the optimal starting point of the prediction is located.
Since the M4 data cover a variety of frequency data such as monthly, quarterly, and annual data, and there are tens of thousands of each frequency data, the model constructed in this paper can more adequately identify the intervals where the corresponding optimal starting points are located based on various types of time series characteristics, and make effective forecasts.
3.4 Prediction of Optimal Starting point
After training the Optimal Starting Point (OSP) model, we can utilize it to enhance the predictions of the basic model $M$. For a given set of data to be predicted, our improvement method is as follows:
  1. [1)]
  2. 1.
    Extract the features of the data to be predicted and input them into the OSP model to obtain each The optimal starting interval of each time series to be predicted $y_{bst}$
  3. 2.
    Based on the optimal starting interval of prediction, determine the interval position of the data to be predicted. Then, select $n$ points in the middle partition of the interval and use them as starting points to construct $n$ new sequences.
  4. 3.
    Input the $n$ new sequences into the basic prediction model $M$ to obtain $n$ prediction sequences. The final prediction result is then obtained by averaging these n prediction results.
3.5 Evaluation Metrics
In order to maximize prediction accuracy, the objective function of the OPS-TSP model is defined by
$\operatorname{argmin}\ \operatorname{Metric}(\widehat{y},y,m,n),$
where $m$ is the number of interval of each series data, $n$ is the number of point in each sub-interval.
We select evaluation metrics, including MASE, MAPE and calculate the average of all new time series prediction errors and compare them with the original ones.
3.5.1 MASE
MASE(Mean Absolute Scaled Error) is a scale-free error metric that gives each error as a ratio compared to a baseline’s average error:
$\operatorname{MASE}=mean(|q_{t}|),$
where:
$q_{t}=\frac{e_{t}}{\frac{1}{n-1}\Sigma^{n}_{i=2}|y_{i}-y_{i-1}|}.$
3.5.2 MAPE
MAPE(Mean Absolute Percentage Error) is 0% for a perfect model and greater than 100% for an inferior model:
$\operatorname{MAPE}=\frac{100\%}{n}\sum_{i=1}^{n}\Big{|}\frac{\widehat{y_{i}}-y_{i}}{y_{i}}\Big{|}.$
4 Empirical analysis
In this section, we present several empirical analysis to demonstrate the finite sample performances of the proposed OSP-TSP algorithm. Code for reproducing the results is available at https://github.com/feng-li/forecasting-with-optimal-starting-point.
4.1 The M4 dataset
We have chosen the M4 dataset to validate the effectiveness of our proposed Optimal Starting Point (OSP) model construction process. The M4 dataset, compiled by the National Technical University of Athens (NTUA), comprises a total of 100,000 time series collected from several publicly accessible websites, ensuring a diverse and rich set of data sources. Furthermore, the M4 dataset spans multiple industry sectors, including services, tourism, imports and exports, demographics, education, government, bonds, stocks, insurance, loans, real estate, transportation, and natural resources and environment Spiliotis et al. (2020). This wide coverage lends strong representation and applicability to the field of time series forecasting.The specific data types in various fields are shown in Table 2
In addition, the M4 dataset boasts a large capacity, encompassing six data frequencies (annual, quarterly, monthly, weekly, daily, and hourly) and six application areas (micro, macro, financial, demographic, and others). It’s worth noting that, on average, the sequences in the M4 dataset are longer than those in the M3 dataset. This provides more opportunities for complex methods that require substantial amounts of data for effective training Kang et al. (2017). The forecast steps are determined based on the nature of the decisions that the data frequency is most likely to support within a company or organization. For instance, annual data is typically used to support long-term strategic decisions spanning one to five years into the future. Conversely, high-frequency data is typically used to support short-term operational decisions, ranging from a few hours to a few weeks. Makridakis et al. (2020).
Table 2: Number of M4 series per data frequency and domain.
Data Frequency Micro Industry Macro Finance Demographic Other Total
Yearly 6,538 3,716 3,903 6,519 1,088 1,236 23,000
Quarterly 6,020 4,637 5,315 5,305 1,858 865 24,000
Monthly 10,975 10,017 10,016 10,987 5,728 277 48,000
Weekly 112 6 41 164 24 12 359
Daily 1,476 422 127 1,559 10 633 4,227
Hourly 0 0 0 0 0 414 414
Total 25,121 18,798 19,402 24,534 8,708 3,437 100,000
Initially, we select three types of seasonal data (annual, quarterly, and monthly) that have the largest volume for empirical research. For each type of data, 70% of the data is selected as the training set to train the Optimal Starting Point (OSP) prediction model. In this context, the ETS and thetaf models are used as the basic prediction model $M$. The sequence segmentation parameters $m$ and $n$ are set to 5 and 4, respectively, and MASE is used as the minimum target error to train the optimal starting section. The trained OSP model is then used to predict the optimal starting points for the remaining data. These predicted optimal starting points are used to reconstruct the new data and make predictions using ETS and thetaf. We then output the average error of the improved prediction of the reconstructed sequence using the optimal starting point in the test data and compare it with the average error of the prediction using the full data.
Table 3 displays the prediction errors MAPE and MASE on the test set corresponding to the eight optimal starting point prediction methods (the combination of the three conditions mentioned in Chapter 1). MASE, as our training target, directly reflects the prediction improvement, while MAPE error indicates whether other non-target errors can be concurrently improved in this process. From Table 3, we observe that the MASE error predicted by almost all the data modified with the optimal starting point is smaller than that predicted by the complete sequence, indicating that the error improvement effect has been achieved. Regarding the improved results of MAPE error, we note that Improved prediction accuracy is also achieved in most cases.
Table 3: OSP improved predictions for ETS and thetaf models on M4 data.
Model Class Method Yearly Quarterly Monthly Mean
MAPE MASE MAPE MASE MAPE MASE MAPE MASE
ETS Actual xgbcls 17.70 2.67 11.95 0.99 16.69 0.87 15.74 1.33
xgbreg 17.08 2.66 11.95 0.98 16.66 0.87 15.57 1.33
lgbcls 17.62 2.61 12.06 0.99 16.67 0.87 15.74 1.32
lgbreg 16.86 2.68 11.88 0.98 16.72 0.87 15.53 1.34
Average xgbcls 16.49 2.52 11.75 0.95 16.49 0.83 15.29 1.27
xgbreg 16.65 2.67 11.83 0.98 16.72 0.87 15.47 1.33
lgbcls 16.39 2.51 11.83 0.94 16.52 0.83 15.30 1.27
lgbreg 16.58 2.71 11.87 0.98 16.69 0.87 15.45 1.34
Total series 18.06 3.44 12.20 1.15 16.83 0.95 15.96 1.61
thetaf Actual xgbcls 17.20 2.48 12.20 1.04 15.76 0.89 15.21 1.31
xgbreg 16.95 2.51 11.94 1.00 15.70 0.88 15.05 1.30
lgbcls 17.08 2.48 12.27 1.06 15.81 0.89 15.22 1.31
lgbreg 17.01 2.52 11.98 1.00 15.72 0.88 15.08 1.31
Average xgbcls 17.20 2.42 11.89 0.98 15.49 0.85 14.99 1.26
xgbreg 16.87 2.52 11.85 1.00 15.53 0.87 14.93 1.31
lgbcls 16.98 2.47 11.92 0.98 15.50 0.84 14.95 1.27
lgbreg 16.93 2.52 11.87 1.00 15.60 0.88 14.98 1.31
Total series 17.06 3.40 11.82 1.22 15.68 0.97 15.04 1.62
If we were to construct eight different improvement methods to improve the forecast for the same data, it would render the program overly complex. Therefore, we consider opting for only a subset of these improvement methods. The ‘mean’ column in the aforementioned table represents the result of training with the minimum average error as the target. It can be observed that training with the average minimum error as the objective consistently yields better results than training with the absolute minimum. This can be attributed to two factors. Firstly, the instability that arises from intercepting the change of prediction error at the starting point of the time series. Secondly, our final prediction is constructed by taking the four equally divided points of the optimal starting interval of the prediction, and then averaging the values after prediction. This approach is also similar to the idea of minimum average error, which requires the overall prediction results to be better within an interval. Consequently, we can attempt to use only the lowest average error as a prediction.
Acknowledging that the subinterval division may be coarse, we divided each sequence into an average of 10 subintervals ($m=10$), also setting $n=4$, and subsequently output the forecast improvement of our model across the three frequencies. The error results are presented in Table 4. It can be observed that the forecast results, when using the optimal prediction starting point under this model setting, are still significantly superior to those obtained using the complete dataset. Concurrently, we compared the forecast results of the model with those set to be divided into 5 sub-intervals. We found that the model with the finer molecular interval did not yield a significant improvement effect. This suggests that employing more sub-intervals may not necessarily lead to better results, and that the use of fewer sub-intervals can also yield good prediction effects and reduce our workload in the model training phase.
Table 4: Prediction results of OSP model at $m=10$.
Model Class Method Yearly Quarterly Monthly Mean
MAPE MASE MAPE MASE MAPE MASE MAPE MASE
ETS Actual xgbcls 17.18 2.67 11.86 1.00 16.75 0.90 15.62 1.35
xgbreg 17.22 2.70 11.96 0.98 16.87 0.91 15.71 1.36
lgbcls 17.32 2.68 11.88 1.00 16.75 0.91 15.66 1.36
lgbreg 17.00 2.70 11.93 0.98 16.89 0.91 15.66 1.36
Average xgbcls 16.94 2.63 11.96 0.96 16.54 0.88 15.48 1.32
xgbreg 16.80 2.67 11.93 0.97 16.92 0.91 15.63 1.35
lgbcls 16.84 2.57 11.92 0.95 16.52 0.88 15.43 1.31
lgbreg 16.88 2.67 11.99 0.98 16.92 0.91 15.66 1.35
Total series 18.06 3.44 12.20 1.15 16.83 0.95 15.96 1.61
thetaf Actual xgbcls 17.11 2.63 12.35 1.07 15.88 0.92 15.29 1.37
xgbreg 17.12 2.65 12.06 1.00 15.71 0.93 15.13 1.36
lgbcls 17.57 2.62 12.58 1.08 15.91 0.94 15.47 1.38
lgbreg 17.12 2.65 12.03 1.00 15.72 0.93 15.13 1.36
Average xgbcls 16.84 2.65 12.00 1.00 15.95 0.90 15.16 1.35
xgbreg 16.93 2.67 11.92 0.99 15.68 0.92 15.03 1.36
lgbcls 16.91 2.65 12.00 1.00 15.92 0.90 15.17 1.35
lgbreg 16.79 2.67 11.97 1.00 15.71 0.93 15.03 1.37
Total series 17.06 3.40 11.82 1.22 15.68 0.97 15.04 1.62
In order to demonstrate the universality of our improvement method across different frequency data, we continue to evaluate the improved error forecast results based on ETS and thetaf for weekly, daily, and hourly data in the M4 dataset. This time, we only use the method of marking the interval with the minimum average error to train the prediction results after training the model. The results are presented in Table 5. It can be observed that for the improved target MASE, the improved results of all models exhibit higher prediction accuracy than those obtained using the complete dataset. Concurrently, MAPE also shows a certain degree of improvement in some models.
Simultaneously, we have distilled and presented the five most pivotal variables within the framework of various seasonal models, as detailed in Table 6. Notably, regardless of the seasonal model in question, Curvature and Linearity occupy paramount positions, forming the cornerstone features for predicting optimal starting points. For datasets exhibiting pronounced seasonality, $seas\_acf1$ (seasonal autocorrelation at lag 1) and $seasonal\_strength$ consistently rank high across quarterly, monthly, and even hourly models, emphatically demonstrating the indispensable nature of seasonal factors across these diverse timescales.
Furthermore, we observe that the autocorrelation of time series, as another crucial characteristic, manifests itself in differing forms contingent upon the frequency of the data. In low-frequency annual data, the autocorrelation coefficients of the raw series, particularly $x\_acf1$ (autocorrelation at lag 1) and $x\_acf10$ (sum of squares of the first ten autocorrelation coefficients), exhibit heightened significance. However, as the data frequency escalates to weekly, daily, and even hourly levels, the autocorrelation features of differenced series, such as $diff1\_acf1$ (autocorrelation at lag 1 after first-order differencing) and $diff2\_acf1$ (autocorrelation at lag 1 after second-order differencing), become increasingly pivotal. In summary, for more granular time series data, the autocorrelation features of series after multiple differencing emerge as the decisive factors in identifying and determining optimal starting points.
Table 5: Error results for weekly, daily, and hourly data in M4 data.
Model Method Weekly Daily Hourly
MAPE MASE MAPE MASE MAPE MASE
ETS xgbcls 9.17 1.98 4.90 2.71 22.15 1.79
xgbreg 9.21 2.00 5.11 2.77 21.65 1.61
lgbcls 9.22 2.00 4.95 2.69 21.68 1.72
lgbreg 9.11 1.99 5.08 2.77 22.09 1.71
Total series 9.17 2.50 4.81 3.28 23.15 1.99
thetaf xgbcls 9.19 2.09 4.97 2.70 25.04 1.63
xgbreg 9.25 2.12 5.35 2.77 24.52 1.62
lgbcls 9.19 2.04 5.12 2.72 24.34 1.61
lgbreg 9.27 2.12 5.42 2.79 24.50 1.64
Total series 9.05 2.66 5.58 3.25 23.57 2.60
Table 6: The importance of the top 5 features under different seasonal models.
Order Yearly Quarterly Monthly Weekly Daily Hourly
1 curvature curvature linearity curvature curvature seas_acf1
2 linearity linearity length linearity linearity curvature
3 x_acf1 length curvature diff1_acf1 diff1_acf10 linearity
4 trend seas_acf1 seas_acf1 x_acf1 spike diff2_acf1
5 x_acf10 seasonal_strength seasonal_strength diff1_acf10 length diff2_acf10
4.2 Real world data
Given that the M4 dataset is an artificially constructed dataset, and although we know it originates from various industries in reality, it’s challenging for us to identify its exact source. Therefore, we resort to using data from more transparent sources for empirical study. We selected some data for testing, and the experiment was conducted on five real-world datasets - GDP, Construction Industry, Exchange Rate, Confidence Index, and Import Value, detailed in the appendix. We also employ metrics such as Mean Absolute Scaling Error (MASE) and Mean Absolute Percentage Error (MAPE) to evaluate the performance of our proposed approach. Each dataset is partitioned into a training set and a test set, adhering to a 70/30 split, where 70% of the data is allocated for training and the remaining 30% is reserved for testing. We train the model on the training set, and then use the trained model on the remaining dataset to predict the optimal starting point. Subsequently, we construct a new time series to improve the prediction results. The prediction results on the dataset are output as follows. It can be observed that improvements on these datasets using the optimal starting point of prediction of our model consistently yield better prediction results. Meanwhile, to reflect the adaptability of our model’s prediction, it can theoretically be applied to any basic prediction method. Therefore, we supplement the prediction results of ARIMA and nnetar-based prediction models, and we can observe that our use of the optimal starting point can also achieve better prediction results in these two models.
Table 7: The prediction results of the real dataset via OSP under four basic prediction methods.
Model Method Yearly Quarterly Monthly
GDP Industry Exchange rate Confidence index Amount of imports
MAPE MASE MAPE MASE MAPE MASE MAPE MASE MAPE MASE
ETS xgbcls 13.89 3.48 9.00 1.08 10.71 1.57 2.08 1.56 22.39 1.59
xgbreg 13.66 4.06 9.90 1.12 9.36 1.23 2.13 1.29 26.04 1.90
lgbcls 13.72 3.53 10.18 1.18 10.74 1.55 2.24 1.83 26.06 1.91
lgbreg 13.55 3.96 10.18 1.18 10.74 1.41 2.18 1.18 26.09 1.9
Total series 13.97 4.57 11.21 1.42 10.93 1.43 1.72 1.48 24.37 1.61
thetaf xgbreg 12.58 3.64 12.26 1.40 8.25 1.15 1.24 0.68 23.1 1.66
lgbcls 13.01 3.46 12.12 1.40 7.48 1.08 1.42 1.13 25.16 1.79
lgbreg 12.48 3.67 12.12 1.40 7.98 1.16 1.42 0.76 25.16 1.79
xgbcls 12.23 3.34 12.17 1.41 8.25 1.19 1.14 0.74 23.05 1.65
Total series 12.99 4.73 14.89 1.7 7.29 0.95 1.14 1.01 20.65 1.52
ARIMA xgbcls 13.19 3.28 19.33 1.51 10.6 1.55 1.91 1.52 23.58 0.95
xgbreg 14.22 3.68 19.25 1.40 10.66 1.46 1.95 1.07 22.05 0.87
lgbcls 13.51 3.3 18.39 1.40 9.98 1.5 1.78 1.48 23.53 1.08
lgbreg 16.49 4.55 19.22 1.34 9.98 1.33 1.78 0.98 21.94 0.86
Total series 12.09 3.93 18.96 1.73 9.09 1.2 1.21 1.08 29.1 1.97
nnetar xgbcls 14.92 3.84 22.18 1.82 9.03 1.44 1.89 1.18 36.16 1.41
xgbreg 14.1 3.99 27.94 2.47 8.4 1.29 1.24 0.65 30.21 1.24
lgbcls 14.34 3.69 28.31 2.22 9.23 1.51 1.03 0.68 29.13 1.37
lgbreg 14.73 4.17 27.22 2.64 8.89 1.46 0.82 0.42 36.88 1.24
Total series 15.22 5.11 26.23 2.59 9.57 1.31 0.98 0.84 42.29 3.03
However, it’s important to consider that we may not always have access to large datasets, such as the M4 dataset, for training. If we only have a small amount of data, directly extracting features for model training may not yield satisfactory results. To address this situation, we propose two solutions. The first approach involves using a previously trained model. If we have trained the OSP model on another large dataset, we can directly use it to make predictions on the current dataset. Although we advocate for the training data to be similar to the prediction dataset, considering the large volume of data used to train the model in the previous dataset, it can learn to a certain extent to generate corresponding predictions for similar situations in the current dataset’s features. The second method takes into account that our model has relatively flexible requirements for training datasets. We can consider using a generated simulation dataset for training. For instance, GRATIS is capable of efficiently generating new time series with controllable features Kang et al. (2019). This method can effectively augment our original dataset. We can train the model on the newly generated data and then apply it to the actual data to enhance the prediction, a strategy that has proven to be effective.
Table 8: The prediction results of the OSP model via trained M4 data.
Model Method Yearly Quarterly Monthly
GDP Industry Exchange rate Confidence index Amount of imports
MAPE MASE MAPE MASE MAPE MASE MAPE MASE MAPE MASE
ETS xgbcls 14.85 2.61 9.45 0.94 11.14 1.68 2.05 1.52 28.27 1.10
xgbreg 15.09 2.32 9.00 0.91 10.71 1.55 2.06 1.50 28.60 1.02
lgbcls 15.41 2.72 9.22 0.91 12.47 1.93 2.11 1.54 27.50 1.13
lgbreg 15.11 2.36 9.16 0.94 16.31 2.23 2.10 1.58 27.58 1.04
Total series 16.74 5.13 11.21 1.42 11.92 1.74 2.00 1.57 27.13 1.60
thetaf xgbcls 12.40 2.23 11.83 1.14 9.34 1.46 1.53 1.07 28.24 1.13
xgbreg 12.54 2.06 11.92 1.16 9.35 1.26 1.54 1.17 28.26 1.08
lgbcls 13.29 2.36 12.24 1.16 9.30 1.46 1.65 1.19 30.22 1.19
lgbreg 12.88 2.13 11.99 1.17 9.29 1.25 1.58 1.20 28.92 1.12
Total series 14.20 4.99 14.89 1.70 8.99 1.35 1.56 1.19 26.89 1.63
Table 9: The prediction results of the OSP model via simulated data from GRATIS.
Model Method Yearly Quarterly Monthly
GDP Industry Exchange rate Confidence index Amount of imports
MAPE MASE MAPE MASE MAPE MASE MAPE MASE MAPE MASE
ETS xgbcls 15.14 3.03 9.65 0.99 11.62 1.82 2.11 1.58 26.05 1.18
xgbreg 14.95 2.67 10.91 1.02 10.93 1.61 2.04 1.47 27.70 1.19
lgbcls 14.96 2.96 10.49 1.11 11.62 1.72 2.12 1.54 26.10 1.20
lgbreg 14.67 2.60 9.54 1.05 11.77 1.72 2.08 1.48 26.39 1.16
Total series 16.74 5.13 11.21 1.42 11.92 1.74 2.00 1.57 27.13 1.60
thetaf xgbcls 13.96 2.87 12.03 1.12 10.09 1.48 1.72 1.21 27.62 1.16
xgbreg 13.17 2.48 12.25 1.20 9.08 1.38 1.54 1.01 27.56 1.10
lgbcls 13.45 2.81 12.15 1.17 9.28 1.35 1.61 1.14 27.86 1.16
lgbreg 12.77 2.33 12.25 1.21 9.22 1.39 1.58 0.97 27.53 1.08
Total series 14.20 4.99 14.89 1.70 8.99 1.35 1.56 1.19 26.89 1.63
4.3 Discussion of other content
Forecast combination In the empirical section, we consistently used the interval with the smallest average error as the target for training the model, ultimately generating prediction results for four distinct settings. However, when we have no prior knowledge about the future, choosing one out of the four prediction results can be challenging. Therefore, we suggest considering the method of forecast combination. The most straightforward approach is to average all the forecast results and use them as the final forecast result. Alternatively, we can average the improved prediction results of the classification model or regression model separately and then use them as predictions.
Additionally, we propose that an auxiliary meta-model can be trained to pre-assign weights to each prediction result to obtain the final combined prediction. For instance, when training the meta-model, you can input the features of the training set data, with the objective being to minimize the errors of the four prediction results of the training set data. This allows it to assign different prediction weights to new time series data.
In the table, we output the combined forecast results using the average method in the M4 data forecast results. It can be seen that although the combined forecast results are not as good as the previous optimal forecast, they can still be effectively improved compared to using the total time series. This also resolves the difficulty of not knowing how to choose the final prediction.
Table 10: Prediction error of combination prediction and random starting point.
Model Method Yearly Quarterly Monthly
MAPE MASE MAPE MASE MAPE MASE
ETS Total series 18.06 3.44 12.20 1.15 16.83 0.95
Full model 16.19 2.62 11.67 0.97 16.46 0.85
classification 16.27 2.51 11.72 0.94 16.45 0.83
Regression 16.53 2.70 11.82 0.98 16.66 0.87
Random 17.11 3.04 11.98 1.10 16.94 1.00
thetaf Total series 17.06 3.40 11.82 1.22 15.68 0.97
Full model 16.82 2.51 11.79 0.99 15.44 0.86
classification 16.98 2.41 11.86 0.97 15.46 0.84
Regression 16.84 2.53 11.83 1.01 15.54 0.88
Random 17.07 3.01 12.01 1.14 16.72 1.04
Verification of Model Effectiveness To further validate the efficacy of the model and to demonstrate that the improvement in prediction is not merely due to time series interception, we compared the OSP model with a random strategy. In the random strategy, we randomly select an interval as the starting interval, then select ‘n’ equally divided points as the starting points of the new sequence, construct the new sequence, make predictions, and then average the prediction results. The prediction results of the random strategy are presented in the table. It can be observed that our OSP model significantly outperforms the random strategy. Concurrently, the table also indicates that incorrect interception of data for prediction can lead to a decrease in prediction accuracy.
5 Conclusions and future work
This paper presents a novel OSP-TSP approach for optimizing the forecast performance by accurately determining the optimal starting point of the input time series. To simplify the model training, we narrow the prediction goal to focus on predicting the optimal forecast starting interval. Specifically, the training data is evenly divided into multiple sub-intervals, where the prediction errors at equally spaced points within each sub-interval are calculated and recorded. The interval containing the optimal starting point is then identified and marked. We employ XGBoost and LightGBM models to predict this optimal starting interval. In the final prediction step, predictions are generated from a few selected, evenly distributed points within the identified interval. These predictions are used to construct tailored sequences for the baseline model. By averaging the outputs, we achieve a final forecast that reflects enhanced accuracy.
To demonstrate the effectiveness of the proposed method, we conducted several empirical analysis using the M4 dataset as well as several other real-world datasets. The results indicate that forecasts utilizing optimal starting points yield greater accuracy compared to those baseline models which use the complete data. Moreover, with respect to improvement, applying the optimal starting interval significantly reduces target prediction errors while effectively mitigating other types of prediction errors.
Furthermore, we propose two approaches for addressing data scarcity issue. The first approach involves making predictions with a pre-trained model, which can fully leverage the knowledge and experience of the existing model. The second approach uses a generated simulation datasets for training. For example, tools like GRATIS can be employed to generate simulated time series with controllable characteristics for effectively augmenting the original datasets. Models trained on these simulated datasets can then be applied to real-world data so that the prediction accuracy can be effectively enhanced.
We conclude this article with several interesting future topics. First, this study lacks an in-depth analysis of the specific selection criteria for multiple improved results provided by the OSP approach. It is of great interest to train an additional model, which can assign weights to the different improved forecasts for a even better prediction performance. Moreover, this paper only conducts empirical research using ETS and thetaf as the baseline models. In theory, the choice of baseline models can be more diverse and advanced. Future work could involve a more comprehensive empirical analysis along this direction, exploring a wider range of baseline models to discover additional opportunities for enhancing prediction accuracy.
Acknowledgements
This research is supported by the National Social Science Fund of China (22BTJ028)
[…]